OpenStudy (anonymous):

6 person are seated at a round table. What is the probability that 2 particular persons sit together?

OpenStudy (perl):

call the two particular persons, A, B

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@ganeshie8 . Number of ways 6 persons sit on a round table = (6 - 1)! = 5! .-. n(S) = 5! Since two persons always sit together, we con- sider them one. So, total persons are 4 + 1 = 5 .•. Number of ways these 5 persons sits on the round table is (5 - 1)! = 4! ways and these two persons sit in 2! ways among themselves. N(E)=4!*2! N(E)/N(S)=2/5

OpenStudy (mayankdevnani):

put A and B in the box `AB`

OpenStudy (mayankdevnani):

now total persons :- `AB`P1,P2,P3,P4 total number of persons=5

OpenStudy (mayankdevnani):

arrange them in a circle :- (n-1) !=((5-1)!

OpenStudy (mayankdevnani):

total outcomes :- (6-1) !

OpenStudy (mayankdevnani):

and we can arrange AB in two ways like AB and BA so,multiply by 2 ! so,we get 4 !*2!

OpenStudy (mayankdevnani):

\[\large \bf probability=\frac{4! \times 2!}{5!}\]

OpenStudy (anonymous):

why we taking AB in one box.that is confusing meand why we are taking circular seating for (n-1)!

OpenStudy (mayankdevnani):

i mean the two particular persons be A and B and they want to sit together so i put them in box or simply you can say that they never separate

OpenStudy (mayankdevnani):

REMEMBER :- In circular permutation,we arrange `n` people in `(n-1) !` and in linear or simple permutation,we arrange `n`people in `n!`

OpenStudy (mayankdevnani):

hope you understand @dg2